A family comes home from a long vacation with laundry to do and showers to take. The water heater has been turned off during vacation. If the water heater has a capacity of 48.8 gallons and a 4890 W heating element, how much time is required to raise the temperature of the water from 21.1°C to 56.9°C? Assume that the heater is well insulated and no water is withdrawn from the tank during this time.
Answer #1

m*cp*dt = p*t

m = 21.1gal(0.18m3)*998kg/m3 = 184.36

184.36 * 4,187 * 35.8 = 4.910 * t

t= 5628 sec